Test the series: infinity sigma n=1 (-1)^ n-1 sin nx n sqrt n for absolute and conditional convergence

1
Expert's answer
2022-01-17T15:26:46-0500

"\sum (-1)^ {n-1}\frac{ sin nx}{ n \sqrt {n}}"


"a_n=|(-1)^ {n-1}\frac{ sin nx}{ n \sqrt {n}}|=\frac{ |sin nx|}{ n \sqrt {n}}"


"b_n=\frac{1}{n\sqrt n}"


since "a_n\le b_n" and "\sum b_n" converges , "\sum a_n" converges as well


so, series "\sum (-1)^ {n-1}\frac{ sin nx}{ n \sqrt {n}}" converges absolutely